# More than Meets the Eye: Understanding the Effect of Individual Objects on Perceived Visual Privacy

### v1

@@ 5.1. Limitations and Future Work
In this section, we discuss the limitations of our work in terms of methodological trade-offs and factors affecting generalizability, which may constrain the external validity of our results. We also outline directions for future research to address these limitations.
@@
Most of the participants in our user study identify as White/Caucasian, highly educated, younger, and mid-career adults. Such a skewed demographic sample could limit the generalizability of our results for cultural and personal background diversity, as these factors tend to shape users’ privacy awareness and perceptions. Similarly, our study involved a representative set of privacy-sensitive objects. Although these represent a good range of sensitive objects that are widely available in visual content, such a finite set may not fully capture the diversity and combinations of objects encountered in the wild. Due to this, our findings may apply to real-world scenarios only to a certain extent. Future work should include more diverse participants and broaden the scope of objects for the user study to capture and understand the complexity of everyday contexts better.
@@
From a methodological point of view, we relied mainly on synthetically generated images to eliminate possible other effects on the perceived privacy level. Although it is a good use case for generative AI in HCI research, and we instructed the participants to imagine themselves as the subject or the photographer/owner, artificial images may not evoke the same kind of cognitive responses as authentic photos or those of the participants themselves. On the other hand, we studied the influence on privacy perception of at most two objects. Real-world visual contents have higher-order combinations, and studying them could open up opportunities to make more detailed observations about co-presence and dominance effects. Furthermore, we relied on conditional branching logic in our user study to personalize the questionnaire and avoid survey fatigue for the participants. Some branches received very few participants and responses, thus limiting statistical comparisons for specific contexts. Future work may consider complementing synthetic stimuli with carefully curated real-world images (with appropriate ethical considerations) and increasing the sample size to ensure sufficient coverage of all conditional branches to validate our results.
6. Conclusion
@@
This paper advances the study of visual privacy by moving beyond a whole-image approach to a fine-grained, object-level perspective. Using a mixed-methods study with 92 participants, we uncovered how specific objects, their co-presence, and contextual cues shape people’s privacy perceptions. Our results highlight the hidden dynamics that guide privacy judgments and demonstrate that small contextual details can substantially influence perceived privacy concerns. We reemphasize the importance of human-centric design approaches that can help simplify the cognitive and technical complexity of visual privacy. Towards this end, we believe our work lays the groundwork for developing adaptive, context-aware systems by providing empirical and practical insights.
@@ Appendix A Appendix
A.1. Prompts Used for Image Generation
@@
A.2. Full Survey (Café Version)
@@
[1–5] Likert scale
Scale: 1 = Not comfortable at all, 2 = Slightly uncomfortable, 3 = Neutral, 4 = Slightly comfortable, 5 = Very comfortable
@@
[1–5] Likert scale
[1–5] Likert scale
[1–5] Likert scale
[1–5] Likert scale
[1–5] Likert scale
Scale: 1 = Not comfortable at all, 2 = Slightly uncomfortable, 3 = Neutral, 4 = Slightly comfortable, 5 = Very comfortable
@@
[1–5] Likert scale
[1–5] Likert scale
[1–5] Likert scale
[1–5] Likert scale
Scale: 1 = Not comfortable at all, 2 = Slightly uncomfortable, 3 = Neutral, 4 = Slightly comfortable, 5 = Very comfortable
@@
A.3. Additional Tables
Table 10. Mapping between comfortability levels and perceived privacy levels.
Comfortability Level Perceived Privacy Level
1 (Not comfortable at all) 5 (Most sensitive)
2 (Slightly uncomfortable) 4
3 (Neutral) 3
4 (Slightly comfortable) 2
5 (Very comfortable) 1 (Least sensitive)
Table 11. Significant demographic effects on comfortability levels. Mann–Whitney U tests were applied for gender, showing that female participants reported higher comfortability levels in all significant cases. Kruskal–Wallis H tests were used for the remaining demographic variables, indicating at least one group distribution differed significantly from the others.
Foreground PSO Background PSO Category $p$-value
Café
Screen E-mail Gender (F >> M) 0.0489
Screen Phone Number Gender (F >> M) 0.0289
Face None Social Usage 0.0332
Paper Face (p.) Social Usage 0.0126
Screen None Social Usage 0.0391
Screen Face (r.) Social Usage 0.0401
Face Landmark Profession 0.0377
Screen None Profession 0.0280
Screen Face (r.) Profession 0.0430
Face Landmark Age 0.0206
Paper Full Name Age 0.0355
Screen E-mail Age 0.0177
Screen Face (r.) Age 0.0206
Office
Face Face (a.) Ethnicity 0.0399
Paper Full Name Ethnicity 0.0056
Paper Signature Ethnicity 0.0391
Screen E-mail Ethnicity 0.0012
Paper Face (p.) Education 0.0427
Face Tattoo Age 0.0393
Screen Date Age 0.0408

### v2

@@ V-A Limitations and Future Work
This subsection discusses the limitations of our work in terms of methodological trade-offs and factors affecting generalizability, which may constrain the external validity of our results, as well as outline future research directions.
@@
Most of the participants in our study identify themselves as White/Caucasian, highly educated, younger and mid-career adults. This skewed demographic sample may limit the generalizability of our results regarding cultural and personal background diversity, as such factors shape users’ privacy awareness and perceptions. Likewise, although we used a representative set of privacy-sensitive objects, this finite set may not capture the full diversity or combinations of objects encountered in the wild, and our findings may apply to real-world scenarios only to a certain extent. Future work should include more diverse participants and broaden the scope of objects to capture the complexity of everyday contexts better.
@@
Methodologically, we relied mainly on synthetically generated images to eliminate possible external effects on perceived privacy. While generative AI is useful in HCI research and we instructed participants to imagine themselves as the subject or photographer, artificial images may not evoke the same cognitive responses as real ones. On the other hand, we studied privacy perception with at most two objects, whereas real-world visual content contains higher-order combinations; examining these could yield more detailed observations about co-presence and dominance effects. Furthermore, conditional branching reduced survey fatigue but resulted in too few responses in some branches, limiting statistical comparisons. Future work may complement synthetic stimuli with ethically curated real-world images and increase the sample size, ensuring sufficient coverage across all branches.
VI Conclusion
@@
This paper advances the study of visual privacy by moving beyond a whole-image approach to a fine-grained, object-level perspective. Using a mixed-methods study with 92 participants, we uncovered how specific objects, their co-presence, and contextual cues shape people’s privacy perceptions. Our results highlight the hidden dynamics that guide privacy judgments and demonstrate that small contextual details can substantially influence perceived privacy concerns. We reemphasize the importance of human-centric design approaches that can help simplify the cognitive and technical complexity of visual privacy. Towards this end, we believe that our work lays the groundwork for developing adaptive, context-aware systems by providing empirical and practical insights.
@@ Appendix A Appendix
A-A Prompts Used for Image Generation
@@
A-B Full Survey (Café Version)
@@
Figure 5: Visual accompanying Q12. Scale: 1 = Not comfortable at all, 2 = Slightly uncomfortable, 3 = Neutral, 4 = Slightly comfortable, 5 = Very comfortable
@@
(a) [1–5] Likert scale
(b) [1–5] Likert scale
(c) [1–5] Likert scale
(d) [1–5] Likert scale
(e) [1–5] Likert scale
Figure 6: Visuals accompanying Q13-17. Likert Scale: 1 = Not comfortable at all, 2 = Slightly uncomfortable, 3 = Neutral, 4 = Slightly comfortable, 5 = Very comfortable
@@
(a) [1–5] Likert scale
(b) [1–5] Likert scale
(c) [1–5] Likert scale
(d) [1–5] Likert scale
Figure 7: Visuals accompanying Q20-23. Likert Scale: 1 = Not comfortable at all, 2 = Slightly uncomfortable, 3 = Neutral, 4 = Slightly comfortable, 5 = Very comfortable
@@
A-C Additional Tables and Figures
TABLE IX: Mapping between comfortability levels and perceived privacy levels.
Comfortability Level Perceived Privacy Level 1 (Not comfortable at all) 5 (Most sensitive) 2 (Slightly uncomfortable) 4 3 (Neutral) 3 4 (Slightly comfortable) 2 5 (Very comfortable) 1 (Least sensitive)
TABLE X: Demographic distribution of survey participants (n=92).
Category Response Count (%) Ethnicity White/Caucasian 58 (63%) Asian 20 (22%) Black/African 6 (7%) Mixed/Multi 3 (3%) I don’t wish to disclose 5 (5%) Age 18–24 18 (20%) 25–34 32 (35%) 35–44 9 (10%) 45–54 29 (32%) 55+ 4 (4%) Gender Female 47 (51%) Male 43 (47%) Other 2 (2%) Education High school graduate 4 (4%) Some college 3 (3%) Bachelor’s degree 32 (35%) Master’s degree 29 (32%) Doctorate 24 (26%) Profession Employed full-time 52 (57%) Employed part-time 8 (9%) Student 28 (30%) Retired 2 (2%) Unemployed 2 (2%)
TABLE XI: Significant demographic effects on comfortability levels. Mann–Whitney U tests were applied for gender, showing that female participants reported higher comfortability levels in all significant cases. Kruskal–Wallis H tests were used for the remaining demographic variables, indicating at least one group distribution differed significantly from the others.
Foreground PSO Background PSO Category $p$-value Café Screen E-mail Gender (F >> M) 0.0489 Screen Phone Number Gender (F >> M) 0.0289 Face None Social Usage 0.0332 Paper Face (p.) Social Usage 0.0126 Screen None Social Usage 0.0391 Screen Face (r.) Social Usage 0.0401 Face Landmark Profession 0.0377 Screen None Profession 0.0280 Screen Face (r.) Profession 0.0430 Face Landmark Age 0.0206 Paper Full Name Age 0.0355 Screen E-mail Age 0.0177 Screen Face (r.) Age 0.0206 Office Face Face (a.) Ethnicity 0.0399 Paper Full Name Ethnicity 0.0056 Paper Signature Ethnicity 0.0391 Screen E-mail Ethnicity 0.0012 Paper Face (p.) Education 0.0427 Face Tattoo Age 0.0393 Screen Date Age 0.0408
Figure 8: Violin Plots of significant Mann-Whitney U test results for café and office environments (H2). Each subplot shows comfortability levels (1–5) for a foreground/background PSO pair; width reflects data density, and orange lines mark group mean differences.